help me solve this problem. I have to make a linear equation to solve this problem NHL player got a total of 215 points he obtained 111 more assists than goals. How many goals did the player have that year
Answer by Theo(13342)

let x = number of goals.
let x + 111 = number of assists.
total points is 2x + 111 = 215
2x = 104
x = 52
-----
if i understood the problem correctly.
-----
he got 52 goals
he got 111 more assists than goals which make the number of assists = 163
-----
the total goals plus assists is 52 + 163 = 215
-----